To whom may it concern,
Two boys, A and B are 600 m apart. They walk towards each other at speeds of 35 m per minute and 25 m per minute respectively. After how many minutes will they meet each other ?
Thank you very much,

Found 2 solutions by terbear1963, funmath:
Answer by terbear1963(43)

1st you have to write an algebraic equation for the word problem. 35x + 25x = 600. 35x is the one boy and 25x is the other boy. By solving the equation - you get that x = 10. By plugging 10 into the original problem you get 35(10) + 25(10) = 600. Therefore, your answer is 10 minutes ( x = the time in minutes ) Your math friend, Terry cats7051@yahoo.com

Answer by funmath(2925)

Distance=rate*time, the total of their distances=600 m
They travel the same amount of time represent that by the variable t.
The rate of A=35 m/min, the rate of B=25m/min.
The distance that A travels=35t, the distance that B travels=25t.
Our equation to solve is:
35t+25t=600
60t=600
60t/60=600/60
t=10 min
They travel for 10 minutes before they meet each other.